A body starts from rest and moves with acceleration a0 After t0 it moves with constant velocity for further time t0 and then retards with 2a0 What is the av vel

Grade:11

3 Answers

Arun
3 years ago
Dear student
 
Please tell what is meaning of av
 
I didn't get it.
Please attach clear image if possible
Vikas TU
3 years ago
Dear student 
initial velocity (u)=0
after t0, velocity=a0×t0(case i)
this body travelled next t0 seconds with velocity a0×t0 (case ii)and then retards with 2a0
final velocity=0(case iii)
Case(i)
Displacement=a0(t0)^2/2
time=t0
Case (ii)
Displacement=a0(t0)^2
time=t0
Case (iii)
time=t0/2 [v=u+at]
Displacement=3a0(t0)^2/4
average velocity=total displacement/time
=9a0(t0)^2/4÷5t0/2
=9a0t0/10
Kareena
2 years ago
If it accelerates with a0 for time t0 thereafter retards with 2a0 then the stopping time will be t0/2. The maximum velocity is a0t0. The area will give displacement hence the displacement is [7a0(t0)^2/4] and total time is 5t0/2. Therefore average velocity = displacement /time=7a0t0/10